# Mass gap in non-perturbative quadratic $\mathcal{R}^{2}$ gravity via Dyson-Schwinger

### v1

@@ Abstract
We apply in a simple model derived from quadratic $\mathcal{R}^{2}$ gravity the technique of Dyson-Schwinger equations to solve for its corresponding quantum theory. Particularly, we solve the classical equations of motion to get a solution to the hierarchy of Dyson-Schwinger equations in the limit of large coupling. The model is obtained by looking for conformally flat solutions. The model admits Higgs-like solutions with a single particle having a finite mass. Assuming a Fubini-Lipatov instanton for the vacuum instead, the spectrum is given by an infinite set of excitations of finite mass. In both cases, the model develops a mass gap due to the presence of the quadratic Ricci scalar, after the breaking of the conformal symmetry.
@@ I Introduction
The paper is organised as follows: In Sec. II, we introduce the model and derive the scalar field theory arising from it. In Sec. III, we discuss the classical solutions for the scalar field theory. In Sec. IV, we quantize the model ans solve the set of Dyson-Schwinger equations both with a null or non-null the cosmological constant. In Sec.V, we provide the conclusions.
@@ II Scalar Model for Quadratic Starobinsky gravity: A small review
We present a scalar field model from the full action of the quadratic gravity Starobinsky:1980te ; Mukhanov:1981xt ; Kaneda:2010ut and then we will quantize it with the technique of the Dyson-Schwinger equations.
$S=-\frac{M_{p}^{2}}{2}\int d^{4}x\left(\mathcal{R}-2\Lambda+\frac{\mathcal{R}^{2}}{6M^{2}}\right),$ (1)
@@
$g_{\mu\nu}\rightarrow e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}g_{\mu\nu}.$ (2)
which gives:
$S=\int{\rm d}^{4}x\sqrt{-g}\left[-\frac{M_{\rm p}^{2}}{2}(\mathcal{R}-2\Lambda)+\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi+\frac{3}{4}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}\right].$ (3)
@@
$\mathcal{R}_{\mu\nu}=-M_{p}^{-2}\left[\partial_{\mu}\chi\partial_{\nu}\chi-g_{\mu\nu}\left(\partial_{\sigma}\chi\partial^{\sigma}\chi+\frac{3}{2}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}\right)-2g_{\mu\nu}M_{p}^{2}\Lambda\right]$
$\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}\chi=-\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}.$ (4)
Now after removing the term $\partial_{\mu}\chi\partial_{\nu}\chi$ by the proper gauge choice in harmonic coordinates (see Appendix A) and assuming a Ricci flat metric, as per Minkowski metric, we get the following set of simplified equations:
$\partial_{\sigma}\chi\partial^{\sigma}\chi+\frac{3}{2}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}+2M_{p}^{2}\Lambda=0$
$\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}\chi=-\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}.$ (5)
This yields an overdetermined system of equations, assuming the manifold has a constant determinant as we are considering conformally flat metrics,
$\partial_{\sigma}\chi\partial^{\sigma}\chi+\frac{3}{2}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}+2M_{p}^{2}\Lambda=0$
$\Box\chi=-\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}.$ (6)
For $\chi\gg M_{p}$, the system is consistent yielding a solution:
$\chi_{\infty}=p_{\mu}x^{\mu}+\theta,$ (7)
where $p_{\mu}$ and $\theta$ are the appropriate integration constants, $x^{\mu}$ space-time coordinates. This yields $p^{2}+(3/2)M_{\rm p}^{2}M^{2}+2M_{\rm p}^{2}\Lambda=0$ for the determination of the constants $p_{\mu}$ Anyway, if we assume universe of be isotropic, this solution just recover the well-known (Anti)de Sitter solution in conformal coordinates if we assume $\Lambda<0$ and $|\Lambda|>M^{2}$, otherwise, we can neglect this solution as unphysical. Taking a dimensionless rescaled field variable $\varphi=\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}$, rescaled the dimensionful field $\chi$ with respect to the Planck mass we turn back to the full set of eq.(II), using which we further get:
$\partial_{\sigma}\varphi\partial^{\sigma}\varphi+M^{2}\left(1-e^{-\varphi}\right)^{2}+\frac{4}{3}\Lambda=0$
$\Box\varphi=-M^{2}\left(1-e^{-\varphi}\right)e^{-\varphi},$ (8)
with the Dálembertain operator, $\Box=\partial_{t}^{2}-\Delta_{2}$. Here in this computation $\partial_{t}$ and $\partial_{\sigma}$ denote the partial differentiation with respect to $t$ and $\sigma$ respectively. Let us take $\varphi=-\log Z$, we get
$Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z+M^{2}\left(1-Z\right)^{2}+\frac{4}{3}\Lambda=0$
$-Z^{-1}\Box Z+Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z=-M^{2}\left(1-Z\right)Z.$ (9)
The insert the first equation into to second obtaining
$-Z^{-1}\Box Z-M^{2}\left(1-Z\right)^{2}-\frac{4}{3}\Lambda=-M^{2}(1-Z)Z.$ (10)
$-Z^{-1}\Box Z-M^{2}\left(1-2Z+Z^{2}\right)-\frac{4}{3}\Lambda=-M^{2}(Z-Z^{2}).$ (11)
$-Z^{-1}\Box Z=-M^{2}(3Z-2Z^{2})+M^{2}+\frac{4}{3}\Lambda.$ (12)
$\Box Z=3M^{2}Z^{2}-2M^{2}Z^{3}-M^{2}Z-\frac{4}{3}\Lambda Z.$ (13)
@@
$Z=\frac{1}{\sqrt{2}M}\phi+\frac{1}{2},$ (14)
@@
$\Box\phi=-\frac{2\sqrt{2}}{3}M\Lambda+\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)\phi-\phi^{3}.$ (15)
@@
$V(\phi)=\frac{2\sqrt{2}}{3}M\Lambda\phi-\frac{1}{2}\left(-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}\right)\phi^{2}+\frac{1}{4}\phi^{4}.$ (16)
@@
Figure 1: Plot of the potential of the conformal factor. It is seen the possibility of false vacuum decay for $3M^{2}/16>\Lambda$. All the physical quantity are dimensionless being scaled with the Planck mass $M_{P}$.
Fig. 1 shows different potentials varying the Starobinsky mass $M$ with respect to the Planck mass for a given value of the cosmological constant. The characteristic shape of a false vacuum decay potential can be recognized.
@@ III Classical solutions
III.1 $\Lambda\neq 0$
We can solve eq.(15) when we consider the term $-\frac{2\sqrt{2}}{3}M\Lambda$ as a perturbation. Thus, we write
$\phi(x)=\sum_{n=0}^{\infty}\epsilon^{n}\phi_{n}(x)$ (17)
where we have introduced a fictitious parameter $\epsilon$ for bookkeeping the order. We will set $\epsilon=1$ at the end of computation. We will get
$\sum_{n=0}^{\infty}\epsilon^{n}\Box\phi_{n}(x)=-\epsilon\frac{2\sqrt{2}}{3}M\Lambda+\mu_{0}^{2}\sum_{n=0}^{\infty}\epsilon^{n}\phi_{n}(x)-[\phi_{0}^{3}+3\epsilon\phi_{1}\phi_{0}^{2}+\epsilon^{2}(3\phi_{1}^{2}\phi_{0}+3\phi_{0}^{2}\phi_{2})+\epsilon^{3}(\phi_{1}^{3}+6\phi_{0}\phi_{1}\phi_{2}+3\phi_{0}^{2}\phi_{3})+\ldots].$ (18)
Here $\mu_{0}^{2}=-\frac{4}{3}\Lambda+\frac{1}{2}M^{2}$. This will yield the set of equations
$\Box\phi_{0}$ $=$ $\mu_{0}^{2}\phi_{0}-\phi_{0}^{3},$
$\Box\phi_{1}$ $=$ $\mu_{0}^{2}\phi_{1}-3\phi_{0}^{2}\phi_{1}-\frac{2\sqrt{2}}{3}M\Lambda,$
$\Box\phi_{2}$ $=$ $\mu_{0}^{2}\phi_{2}-(3\phi_{1}^{2}\phi_{0}+3\phi_{0}^{2}\phi_{2}),$
$\Box\phi_{3}$ $=$ $\mu_{0}^{2}\phi_{3}-(\phi_{1}^{3}+6\phi_{0}\phi_{1}\phi_{2}+3\phi_{0}^{2}\phi_{3}),$ (19)
$\vdots$ .
We get the following solutions (this class of solutions arises from our preceding studies for different field theories Frasca:2015yva ; Frasca:2021yuu ; Frasca:2021mhi ; Frasca:2022lwp ; Chaichian:2018cyv ; Frasca:2022vvp ; Frasca:2015wva ; Frasca:2017slg )
$\phi_{0}(x)=\sqrt{\frac{2}{3}}|\mu_{0}|\text{dn}\left(p\cdot x+\theta,-1\right),$ (20)
where dn is a Jacobian elliptic function and having fixed the dispersion relation to
$p^{2}=\frac{1}{3}\mu_{0}^{2},$ (21)
and where $\theta$ is an arbitrary constant arising from integration. To solve for $\phi_{1}$, we need the Green function $C_{1}$ given by the following Lamé equation
$\Box C_{1}(x,x^{\prime})=(\mu_{0}^{2}-3\phi_{0}^{2}(x))C_{1}(x,x^{\prime})+\delta^{4}(x-x^{\prime}).$ (22)
We can get a straightforward solution by noting that the homogeneous equation has the solution
$C_{1h}(x)=\text{cn}\left(p\cdot x+\theta,-1\right)\text{sn}\left(p\cdot x+\theta,-1\right),$ (23)
where cn and sn are Jacobian elliptic functions. Thus, using the results given in Ref. Frasca:2023uaw , we obtain in momentum space
where $K(-1)$ is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{|\mu_{0}|}{\sqrt{3}}.$ (25)
It should be noticed the presence of a zero mass excitations in the spectrum but it is not propagating as $C_{1}(p)$ ha a null contribution in this case. This applies to our discussion for the quantum case. Thus,
$\phi_{1}(x)$ $=$ $-\frac{2\sqrt{2}}{3}M\Lambda\int d^{4}x^{\prime}C_{1}(x-x^{\prime}),$
$\phi_{2}(x)$ $=$ $-3\left(\frac{2\sqrt{2}}{3}M\Lambda\right)^{2}\int d^{4}x^{\prime}C_{1}(x-x^{\prime})\left[\int d^{4}x^{\prime\prime}C_{1}(x^{\prime}-x^{\prime\prime})\right]^{2}\phi_{0}(x^{\prime}),$
$\phi_{3}(x)$ $=$ $-\left(\frac{2\sqrt{2}}{3}M\Lambda\right)^{3}\left[\int d^{4}x^{\prime}C_{1}(x-x^{\prime})\left[\int d^{4}x^{\prime\prime}C_{1}(x^{\prime}-x^{\prime\prime})\right]^{3}\right.$ (26)
$\left.+18\int d^{4}x^{\prime}C_{1}(x-x^{\prime})\phi_{0}(x^{\prime})\int d^{4}x^{\prime\prime}C_{1}(x^{\prime}-x^{\prime\prime})\int d^{4}x^{\prime\prime\prime}C_{1}(x^{\prime\prime}-x^{\prime\prime\prime})\left[\int d^{4}x^{\prime\prime\prime\prime}C_{1}(x^{\prime\prime\prime}-x^{\prime\prime\prime\prime})\right]^{2}\phi_{0}(x^{\prime\prime\prime\prime})\right],$
$\vdots$ .
III.2 $\Lambda=0$
For $\Lambda=0$, we need to introduce an arbitrary source $j$ and compute the functional series
@@ III.2 $\Lambda=0$
By iterating the derivation with respect to $j$ and starting from the equation of motion
$\Box\phi(x)=\frac{M^{2}}{2}\phi(x)-\phi^{3}(x)+j(x),$ (29)
we get
$\Box\phi_{0}(x)$ $=$ $\frac{M^{2}}{2}\phi_{0}(x)-\phi_{0}^{3}(x),$
$\Box C_{1}(x,x_{1})$ $=$ $\left(\frac{M^{2}}{2}-3\phi_{0}^{2}(x)\right)C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $\left(\frac{M^{2}}{2}-3\phi_{0}^{2}(x)\right)C_{2}(x,x_{1},x_{2})-6\phi_{0}(x)C_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $\left(\frac{M^{2}}{2}-3\phi_{0}^{2}(x)\right)C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-6\phi_{0}(x)C_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (30)
@@
We can use the results of the preceding section to solve this set of equations yielding
$\phi_{0}(x)=\frac{1}{\sqrt{3}}M\text{dn}\left(p\cdot x+\theta,-1\right),$ (31)
where one can see M as the Starobinsky mass in the original Lagrangian 1, dn is a Jacobian elliptic function and
$p^{2}=\frac{1}{6}M^{2}.$ (32)
@@
where $K(-1)$ is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{M}{\sqrt{6}}.$ (34)
@@
It is also possible to get a constant solution at leading order
$\phi_{0}=\frac{M}{\sqrt{2}},$ (36)
that is a Higgs solution. This reduces the set of equations to
$\Box C_{1}(x,x_{1})$ $=$ $-M^{2}C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $-M^{2}C_{2}(x,x_{1},x_{2})-\frac{6}{\sqrt{2}}MC_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $-M^{2}C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-\frac{6}{\sqrt{2}}MC_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (37)
@@
$C_{1}(p)=\frac{1}{p^{2}+M^{2}},$ (38)
and we have a single boson of mass $M$ with translation invariance never broken.

### v2

@@ Abstract
We apply in a simple model derived from quadratic $\mathcal{R}^{2}$ gravity the technique of Dyson-Schwinger equations to solve for its corresponding quantum theory. Particularly, we solve the classical equations of motion to get a solution to the hierarchy of Dyson-Schwinger equations in the limit of large Ricci scalar, assumed to be constant and larger than the square of the Starobinsky mass. Moving to the Einstein frame, the model admits Higgs-like solutions with a single particle having a finite mass. We quantize the scalar field showing the appearing of a mass gap through a Higgs-like solution. The presence of the mass gap, that increases with the square root of the Ricci scalar, shows how the effect of the scalar sector at low-energy becomes ineffective, making it relevant only at short distances.
@@ I Introduction
In this work, we start considering generally non-renormalizable theories and obtain some meaningful renormalizable scalar theories to work with, only under some conditions. There exists several such examples, a well-known example being given by the Nambu-Jona-Lasinio model used to describe strong interactions Klevansky:1992qe . This model is characterized by fermion fields with a quartic interaction between them. As such, this model is not renormalizable and the computations depend on a cut-off. When a bosonization procedure is applied Klevansky:1992qe , a $\sigma$-model emerges that is renormalizable, and could represent a description of the breaking of the chiral symmetry in strong interactions. With this principle in mind, we try to have similar approach here showing that, under proper conditions, some effective scalar models could be extracted from quadratic gravity and can be suitably properly studied using the well-known techniques of quantum field theory as the Dyson-Schwinger equations are. However a concrete re-normalizable model of gravity like Stelle gravity or asymptotic safe gravity or non-local gravity (see Ref. Salvio:2018crh for such a discussion) is beyond the scope of the current paper and will be studied in future.
The paper is organized as follows: In Sec. II, we introduce the model and derive the scalar field theory arising from it. In Sec. III, we discuss the classical solutions for the scalar field theory. In Sec. IV, we quantize the model ans solve the set of Dyson-Schwinger equations both with a null or non-null the cosmological constant. In Sec.V, we provide the conclusions.
@@ II Scalar Model for Quadratic Starobinsky gravity: A small review
We present a scalar field model from the full action of the quadratic gravity Starobinsky:1980te ; Mukhanov:1981xt ; Kaneda:2010ut and then we will quantize it with the technique of the Dyson-Schwinger equations [^2] We do not consider here a scenario like Higgs inflation, involving a term like $\xi\phi^{2}\mathcal{R}$. We only restrict ourselves to a minimally coupled scalar field to gravity as comes out from quadratic gravity..
$S=-\frac{M_{p}^{2}}{2}\int d^{4}x\sqrt{-g}\left(\mathcal{R}-2\Lambda+\frac{\mathcal{R}^{2}}{6M^{2}}\right),$ (1)
@@
$g_{\mu\nu}\rightarrow e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}g_{\mu\nu},$ (2)
where $g_{\mu\nu}$ is a generic metric. This gives Salvio:2018crh :
$S=\int{\rm d}^{4}x\sqrt{-g}\left[-\frac{M_{\rm p}^{2}}{2}(\mathcal{R}-2\Lambda e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}})+\frac{1}{2}g^{\mu\nu}\partial_{\mu}\chi\partial_{\nu}\chi+\frac{3}{4}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}\right].$ (3)
@@
$\mathcal{R}_{\mu\nu}=\frac{1}{M_{p}^{2}}\left\{\partial_{\mu}\chi\,\partial_{\nu}\chi-\tfrac{1}{2}g_{\mu\nu}\,g^{\alpha\beta}\partial_{\alpha}\chi\,\partial_{\beta}\chi-g_{\mu\nu}\left[M_{\rm p}^{2}\Lambda\,e^{-2\sqrt{\tfrac{2}{3}}\tfrac{\chi}{M_{\rm p}}}+\tfrac{3}{4}M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\tfrac{2}{3}}\tfrac{\chi}{M_{\rm p}}}\right)^{2}\right]\right\}+\frac{1}{2}g_{\mu\nu}{\cal R},$
$\Box_{g}\chi=\sqrt{\frac{3}{2}}M^{2}M_{p}e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}\right)-\sqrt{\frac{8}{3}}\Lambda M_{p}e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}$ (4)
where $\Box_{g}=-\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}$ is the Laplace-Beltrami operator.
We look for solution with $R$ being a constant. This yield the set of equations for the scalar field
$\partial_{\sigma}\chi\partial^{\sigma}\chi+3M_{\rm p}^{2}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)^{2}+4M_{p}^{2}\Lambda e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}-M_{p}^{2}\mathcal{R}=0,$
$-\frac{1}{\sqrt{-g}}\partial_{\mu}g^{\mu\nu}\sqrt{-g}\partial_{\nu}\chi=\sqrt{\frac{3}{2}}M_{\rm p}M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}\right)e^{-\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}}-\sqrt{\frac{8}{3}}\Lambda M_{p}e^{-2\sqrt{\frac{2}{3}}\frac{\chi}{M_{p}}}.$ (5)
Taking a dimensionless rescaled field variable $\varphi=\sqrt{\frac{2}{3}}\frac{\chi}{M_{\rm p}}$, we further get:
$\partial_{\sigma}\varphi\partial^{\sigma}\varphi+2M^{2}\left(1-e^{-\varphi}\right)^{2}+\frac{8}{3}\Lambda e^{-2\varphi}-\frac{2}{3}\mathcal{R}=0$
$\Box_{g}\varphi=M^{2}\left(1-e^{-\varphi}\right)e^{-\varphi}-\frac{4}{3}\Lambda e^{-2\varphi}.$ (6)
Let us take $\varphi=-\log Z$, we get
$Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z+2M^{2}\left(1-Z\right)^{2}+\frac{8}{3}\Lambda Z^{2}-\frac{2}{3}\mathcal{R}=0$
$-Z^{-1}\Box_{g}Z+Z^{-2}\partial_{\sigma}Z\partial^{\sigma}Z=M^{2}\left(1-Z\right)Z-\frac{4}{3}\Lambda Z^{2}.$ (7)
We insert the first equation into the second obtaining
$\Box_{g}Z=-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)Z+3M^{2}Z^{2}-\left(M^{2}+\frac{4}{3}\Lambda\right)Z^{3}.$ (8)
@@
$Z=\frac{M^{2}}{M^{2}+\frac{4}{3}\Lambda}+\frac{\phi}{\sqrt{M^{2}+\frac{4}{3}\Lambda}},$ (9)
@@
$\Box_{g}\phi=-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)\frac{M^{2}}{\sqrt{M^{2}+\frac{4}{3}\Lambda}}+\frac{2M^{6}}{\left(M^{2}+\frac{4}{3}\Lambda\right)^{\frac{3}{2}}}$ (10)
$+\left[-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)+\frac{3M^{4}}{M^{2}+\frac{4}{3}\Lambda}\right]\phi-\phi^{3}$
@@
$V(\phi)=\left[\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)\frac{M^{2}}{\sqrt{M^{2}+\frac{4}{3}\Lambda}}-\frac{2M^{6}}{\left(M^{2}+\frac{4}{3}\Lambda\right)^{\frac{3}{2}}}\right]\phi-\frac{1}{2}\left[-\left(2M^{2}-\frac{2}{3}\mathcal{R}\right)+\frac{3M^{4}}{M^{2}+\frac{4}{3}\Lambda}\right]\phi^{2}+\frac{1}{4}\phi^{4}.$ (11)
@@
Figure 1: Plot of the potential of the conformal factor. It is seen the possibility of false vacuum decay for $3M^{2}<\mathcal{R}$. All the physical quantity are dimensionless being scaled with the Planck mass $M_{P}$.
Fig. 1 shows different potentials varying the Starobinsky mass $M$ with respect to the Planck mass for a given value of the cosmological constant. The characteristic shape of a false vacuum decay potential can be recognized provided the condition $3M^{2}<\mathcal{R}$ holds.
The condition $3M^{2}<\mathcal{R}$ (strong coupling regime) is essential for our analysis. It grants the reduction of the Laplace-Beltrami operator to that of the flat space, aside from a proper scale changing. Physically, this means that the quadratic term $\mathcal{R}^{2}$ is prevailing with respect to the linear one of the standard Einstein-Hilbert action. For this reason, the quantum behavior in this regime is the main concern of the present study.
We see that the presence of the cosmological constant is not so relevant and so, in the following we set it to 0. This will yield the equation of motion
$\Box_{g}\phi=\frac{2}{3}\mathcal{R}M+\left[\frac{2}{3}\mathcal{R}+M^{2}\right]\phi-\phi^{3},$ (12)
and the potential
$V(\phi)=-\frac{2}{3}\mathcal{R}M\phi-\frac{1}{2}\left[\frac{2}{3}\mathcal{R}+M^{2}\right]\phi^{2}+\frac{\phi^{4}}{4}.$ (13)
@@ III Classical solutions
For our study, we need to introduce an arbitrary source $j$ and compute the functional series
@@
By iterating the derivation with respect to $j$ and starting from the equation of motion (that holds for $3M^{2}<\mathcal{R}$),
$\Box\phi(x)=\Omega+\mu_{0}^{2}\phi(x)-\phi^{3}(x)+j(x),$ (16)
where $\Omega=\frac{2}{3}\mathcal{R}M$ and $\mu_{0}^{2}=\frac{2}{3}\mathcal{R}+M^{2}=\frac{\Omega}{M}+M^{2}$ having set $\Lambda=0$ to make handling the equations simpler, we get
$\Box\phi_{0}(x)$ $=$ $\Omega+\mu_{0}^{2}\phi_{0}(x)-\phi_{0}^{3}(x),$
$\Box C_{1}(x,x_{1})$ $=$ $\left[\mu_{0}^{2}-3\phi_{0}^{2}(x)\right]C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $\left[\mu_{0}^{2}-3\phi_{0}^{2}(x)\right]C_{2}(x,x_{1},x_{2})-6\phi_{0}(x)C_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $\left[\mu_{0}^{2}-3\phi_{0}^{2}(x)\right]C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-6\phi_{0}(x)C_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (17)
@@
The cubic polynomial on the lhs of the leading order equation has three solution as expected $(-M,(3M-\sqrt{3}\sqrt{3M^{2}+8R})/6,(3M+\sqrt{3}\sqrt{3M^{2}+8R})/6)$. The last one is the true vacuum. Thus, if we select it, we can work out a Higgs-like solution and a massive solution already at this stage. We will discuss this point below. On the other hand, we can exploit the strong coupling condition $3M^{2}<\mathcal{R}$ and neglect the constant $\Omega$ with respect to the other terms in the leading order equation. We can also approximate $\mu_{0}^{2}\approx 2R/3$ that can be seen as a leading order approximation [^3] We observe that we are claiming that $RM\ll R\phi_{0}$ that is $\phi_{0}\gg M$ that is consistent with our approximation discussed in the previous section.. This implies that the leading order equation takes the form
$\Box\phi_{0}(x)\approx\mu_{0}^{2}\phi_{0}(x)-\phi_{0}^{3}(x),$ (18)
being $\mu_{0}^{2}$ the approximate value given above. In this way, we are able to get an alternative approximate solution to this set of equations yielding
$\phi_{0}(x)=\sqrt{\frac{2}{3}}\mu_{0}\text{dn}\left(p\cdot x+\theta,-1\right),$ (19)
where one can see M as the Starobinsky mass in the original Lagrangian (1), dn is a Jacobian elliptic function and
$p^{2}=\frac{\mu_{0}^{2}}{3}.$ (20)
@@
where K(1)K(-1) is the complete elliptic integral of the first kind and
$m_{n}=n\frac{\pi}{K(-1)}\frac{\mu_{0}}{\sqrt{3}}.$ (22)
@@
The Higgs-like case can analyzed by taking
$\phi_{0}\approx\mu_{0}.$ (24)
This reduces the set of equations to
$\Box C_{1}(x,x_{1})$ $=$ $-2\mu_{0}^{2}C_{1}(x,x_{1})+\delta^{4}(x-x_{1}),$
$\Box C_{2}(x,x_{1},x_{2})$ $=$ $-2\mu_{0}^{2}C_{2}(x,x_{1},x_{2})-\frac{6}{\sqrt{2}}MC_{1}(x,x_{2})C_{1}(x,x_{1}),$
$\Box C_{3}(x,x_{1},x_{2},x_{3})$ $=$ $-2\mu_{0}^{2}C_{3}(x,x_{1},x_{2},x_{3})-6C_{1}(x,x_{1})C_{1}(x,x_{2})C_{1}(x,x_{3})-\frac{6}{\sqrt{2}}MC_{2}(x,x_{1},x_{2})C_{1}(x,x_{3})$ (25)
@@
$C_{1}(p)=\frac{1}{p^{2}+2\mu_{0}^{2}},$ (26)
and we have a single boson of mass $\sqrt{2}\mu_{0}$ with translation invariance never broken after quantization. We see that the classical theory displays a massive solution in any case, due to the nonlinear potential we obtained from the Starobinsky model.